# Rare and exclusive few-body decays of the Higgs, Z, W bosons, and the top quark

### v3

@@ II.1.3 Non-Relativistic QCD (NRQCD)
The formation of heavy quarkonium in the decay of a massive particle X is a multi-scale problem involving the hard scale given by $m_{\mathrm{X}}$, the heavy-flavour quark mass $m_{\mathrm{Q}}$, the relative momentum of the heavy quark pair $m_{\mathrm{Q}}v$, and the binding energy of the heavy quark pair $m_{\mathrm{Q}}v^{2}$, where $v$ is the typical relative velocity of the heavy quarks inside the meson. Since $m_{\mathrm{X}}\gg mv\gg mv^{2}\gg\Lambda_{\text{QCD}}$, the simple static limit applied for light quarks is not sufficient for these systems. Therefore, for heavy quark-antiquark bound states such as charmonium, bottomonium, and bottom-charm (Bc) mesons, a third class of models is often used where in addition to the usual expansion in powers of $\alpha_{\text{s}}$, the interactions are organized as an expansion in $v$. Since $v^{2}\approx 0.3,0.1$ for charmonium and bottomonium, respectively, the heavy quarks are nonrelativistic and described in a Non-Relativistic QCD (NRQCD) approach [45]. In the NRQCD framework, the decay width of a heavy particle X into a given quarkonium meson M$(n)$ with quantum numbers $n=\,^{2s+1}L_{J}^{[1,8]}$ (corresponding to a state with spin $s$, orbital angular momentum $L$, total angular momentum $J$, in the $[1]$ or $[8]$ colour singlet or octet representations, respectively), $\mathrm{X}\to\mathrm{M}(n)+\mathrm{Y}$, factorizes as
@@
for the singlet charmonium ground states, where $N_{\mathrm{c}}=3$ is the number of colours. The LDMEs for the CS contributions can be obtained from the electromagnetic decays of quarkonia into dileptons or diphotons [46, 47], or from potential models. The NRQCD encapsulates two different mechanisms to describe the evolution of the heavy-flavour quark pair into a quarkonium meson, such as the colour-singlet model (CSM), and the colour-octet model (COM). In the CSM, the $\mathrm{Q}\overline{\mathrm{Q}}$ pairs are produced in CS states at the hard-scattering scale $m_{\mathrm{X}}$, and their quantum numbers are conserved during hadronization. The COM requires the extra radiation of gluons, and its contributions are more suppressed than in inclusive reactions, but has also been considered in a few cases. Beyond the non relativistic description of NRQCD, the relativistic quark model (RQM) [48] has been also employed in a few exclusive decay processes involving heavy quarkonia. The NRQCD calculations of heavy particle decays into quarkonia have been performed at LO, NLO, or next-to-NLO (NNLO) accuracy, resumming in some cases also leading (LL) or next-to-leading (NLL) logarithms.
As an example, the production of a bottom meson $\rm B\equiv(\mathrm{b}\overline{\mathrm{q}})_{n}=B_{\mathrm{u,d}},B_{\mathrm{s}},B_{\mathrm{c}}$ in the heavy particle decay process $\rm X\to\overline{\rm B}+Y$, can be described in a NRQCD-inspired heavy-quark recombination model [49] where the width takes the form
@@
with $(\mathrm{b}\overline{\mathrm{q}})_{n}$ representing the Fock state of the bottom b and accompanying $\rm\overline{q}$ quark, and the factor $\rho[(\mathrm{b}\overline{\mathrm{q}})_{n}\to\overline{\rm B}]$ is a nonperturbative probability for $(\mathrm{b}\overline{\mathrm{q}})_{n}$ to evolve into the $\rm\overline{\rm B}$ meson. In the case of $\rm B_{c}^{\pm}$ production, the nonperturbative transition $\rm(\mathrm{b}\overline{\mathrm{q}})_{n}\to\overline{\rm B}_{c}$ follows a definite velocity power-counting because the relative velocity of the charm and bottom quarks in the rest frame of the bottom meson (like that of the two bottom quarks in the $\Upsilon$ case) are small. This latter assumption is not fulfilled in the $\rm B_{\mathrm{u,d,s}}$ cases, where the Fock state $(\mathrm{b}\overline{\mathrm{q}})_{n}$ contributions to the $\rm\overline{\rm B}$ meson with different quantum numbers (e.g., in colour or angular momentum) are not necessarily suppressed, as aforementioned (although for practical purposes, the only relevant states are $n=\!^{1}S_{0}^{[1]},^{3}S_{1}^{[1]},^{1}S_{0}^{[8]},^{3}S_{1}^{[8]}$ [49]). By using the heavy-quark spin symmetry [45], one can further reduce the nonperturbative transition probabilities $\rho$ from four to two as [26]
@@
where $m_{\mathrm{b}}$ is the bottom quark mass, $m_{\mathrm{{\overline{q}}}}=m_{\rm\bar{d}},m_{\rm\bar{s}}\approx 0.3$ GeV are the constituent light quark masses. The concrete application of the formulas above for the semiexclusive decay of a top quark into a B-meson plus an up-type quark, $\rm t\to\overline{\rm B}_{(s)}^{0}+q$ (discussed in Section VI.3), results in the following widths at LO in the QCD coupling $\alpha_{\mathrm{s}}$ [26]:
@@
with similar (yet longer) expressions available in Ref. [26] for the $\rm t\to\Upsilon+q$ decays.
@@ II.2 Experimental limits
As we will see below, predictions for rare decays of the electroweak and Higgs bosons, and the top quark, have branching fractions in the $10^{-5}$ to $10^{-15}$ range (or even down to $10^{-22}$ for positronium + photon final states). Such tiny branching fractions are very challenging experimentally, and no such decays have yet been observed for any of the particles. To provide an idea of the size of the data samples of W±, Z, H, and t particles discussed in this work, Table 1 collects their total number produced in past, current, and future colliders. The LEP numbers of W and Z bosons are obtained from the LEP-I and LEP-II electroweak summaries [50, 51], and indicate that no rare decay mode with rates below $10^{-7}$ and $10^{-5}$ for the Z and W bosons, respectively, has ever been probed in the clean experimental conditions of an $\mathrm{e}^{+}\mathrm{e}^{-}$ collider. We note, somehow curiously, that the last LEP-II operation, which integrated 2.46 fb-1 over $\sqrt{s}=189$–209 GeV [52], featured a Higgs cross section of $\sigma_{\mathrm{H}}\approx 3$ fb (adding Higgstrahlung and weak-boson-fusion production processes) [53], and therefore LEP-II did produce a few Higgs(125 GeV) boson counts (however, the Higgs searches at the time were optimized for the associated production of a $m_{\mathrm{H}}=115$ GeV scalar boson plus an onshell Z boson, for which the H(125 GeV) signal would not have been visible). For the next $\mathrm{e}^{+}\mathrm{e}^{-}$ machine, we focus on FCC-ee because it is the planned facility with the largest data samples expected to be collected [54]. The FCC-ee numbers in Table 1 cover the time span of the baseline 15-year program with four interaction points and four dedicated runs under consideration: Z-pole, WW threshold, HZ Higgstrahlung, and $\mathrm{t}\overline{\mathrm{t}}$ threshold. The largest data sample of all will be for Z bosons, for which $6\cdot 10^{12}$ particles will be produced. The number of H bosons at FCC-ee includes those produced in the HZ ($1.45\times 10^{6}$) and $\mathrm{t}\overline{\mathrm{t}}$ ($+330$k events) runs, plus via $\rm WW\to H$ at all runs ($+125$k). Although, in much more complicated background conditions than in $\mathrm{e}^{+}\mathrm{e}^{-}$ collisions, the HL-LHC (with 3 ab-1of p-p collisions at $\sqrt{s}=14$ TeV collected per ATLAS/CMS experiment) and the FCC-hh (p-p at $\sqrt{s}=100$ TeV with 30 ab-1) will truly serve as W, Z, Higgs, and top factories. We also list the number of the four massive particles produced at the Tevatron in p-$\mathrm{p}\mbox{-}\overline{\mathrm{p}}$ collisions at $\sqrt{s}=1.96$ TeV, as there exist still a couple of competitive rare decay limits from the CDF experiment. The hadron collider numbers in Table 1 are obtained from the corresponding production cross sections for each particle multiplied by the integrated luminosity quoted (plus the contribution from $\rm\mathrm{t}\overline{\mathrm{t}}\to W^{+}W^{-}+X$ decays for the W± counting). The cross sections at hadron colliders have been either obtained from the existing literature [55] or, when not readily available or not fully up-to-date, have been recomputed at next-to-next-to-leading-order (NNLO) accuracy with mcfm v.8.0 [56] with the NNPDF3.1_NNLO parton distribution functions (PDFs) [57]. Theoretical (scale, PDF) uncertainties are at most 10% and not quoted.
@@
Table 1: Total number of Higgs, Z, and W bosons, and top quarks produced (or expected to be produced) in $\mathrm{e}^{+}\mathrm{e}^{-}$ collisions at LEP and FCC-ee, as well as in p-$\mathrm{p}\mbox{-}\overline{\mathrm{p}}$ at Tevatron, and in p-p collisions at HL-LHC, and FCC-hh. For $\mathrm{e}^{+}\mathrm{e}^{-}$ colliders, the Z-pole, Higgstrahlung and Higgs weak-fusion, and pair production (WW, $\mathrm{t}\overline{\mathrm{t}}$) cross sections (without ISR) are indicated. At LEP and FCC-ee, the numbers of W bosons and top-quarks consider two bosons and two quarks produced per collision around the $\mathrm{e}^{+}\mathrm{e}^{-}\to\rm W^{+}W^{-},\,\mathrm{t}\overline{\mathrm{t}}$ thresholds. For hadron machines, the integrated luminosities and the NNLO production cross section for each particle are indicated. All top quark numbers are for pair production (i.e., the number of events are multiplied by two), and the number of W bosons is the (W+ + W-) sum including also the contributions from $\mathrm{t}\overline{\mathrm{t}}$ decays.
@@
$\sigma$(W) $N$(W) $\sigma$(Z) $N$(Z) $\sigma$(H) $N$(H) $\sigma$($\mathrm{t}\overline{\mathrm{t}}$) $N$(top)
@@
FCC-ee 2–10 pb $5\times 10^{8}$ 59 nb $6\times 10^{12}$ 200, 30 fb $1.9\times 10^{6}$ 0.5 pb $3.8\times 10^{6}$
@@
The largest yield increases across colliders are factors of $6\cdot 10^{3}$, $3\cdot 10^{5}$, and $4\cdot 10^{5}$ for W, Z, and H bosons, respectively, going from LEP to FCC-ee, and a factor of $9\cdot 10^{4}$ for top quarks from Tevatron to HL-LHC. As of today (end of 2024), measurements at LEP, Tevatron, and LHC have been able to set upper limits at 95% confidence level (CL) in about one-fourth of the $\sim$200 rare channels considered here. The most stringent experimental rare decays limits are listed in the tables below, including in some cases new results which are not yet available in the 2024 PDG decays listings [58]. One key goal of our work is to provide reasonable expectations of the achievable upper bounds at the HL-LHC, FCC-ee, and FCC-hh. The corresponding extrapolations are obtained through three different means:
@@
    For a few decay channels, there exist dedicated ATLAS and/or CMS studies that have determined the expected limits at the end of the HL-LHC phase [59, 60, 61, 62, 63, 64]. For such cases, we provide directly the expected limits (divided by $\sqrt{2}$ to account for the statistical combination of two experiments, ATLAS + CMS) with their bibliographical reference.
@@
    For those channels where LHC limits exist today in measurements with a given integrated luminosity (labelled e.g., “$\mathcal{L}_{\mathrm{int}}$(13 TeV)” next), but for which no dedicated studies exist for HL-LHC, we estimate the latter by assuming that they will be statistically improved by the size of the final data sample, namely by the ratio of squared-root integrated luminosities $\sqrt{\smash{2\times 3~{}\mathrm{ab}^{-1}}/\smash{\mathcal{L}_{\mathrm{int}}(\text{13\,TeV})}}$, where the factor of two assumes an ATLAS + CMS combination. For individual LHC measurements carried out with the full Run-2 integrated luminosity of around 140 fb-1, this translates into an expected HL-LHC improvement of more than a factor of six. Bounds estimated that way are conservative, as they ignore improvements in the data analyses (e.g., optimization of the event selection, enlarged categorization depending on the production mode, adoption of more advanced statistical profiling methods, etc.), increased particle production cross sections (from collision energies rising from 13 to 14 TeV), and possible multi-experiment combinations (e.g., adding results from LHCb) for some channels. We have checked that our simple statistical approximation here gives limits comparable to those obtained with the method 1) above whenever dedicated studies are available.
@@
Finally, for the FCC-ee and FCC-hh cases, whenever a given decay channel has a branching fraction commensurate with the expected number of particles produced (i.e., whenever ×$\mathcal{B}\times N_{\mathrm{X}}\gtrsim 1$, where $N_{\mathrm{X}}$ are the FCC numbers given in Table 1 for any given particle), it will appear listed as “producible at FCC-ee/FCC-hh”. Assuming that potential backgrounds are small and/or well under control, as is often the case at $\mathrm{e}^{+}\mathrm{e}^{-}$ colliders, and even more so at FCC-ee where very precise detector performances are being required [54], observation can be achieved with a handful of events. Such an assumption is, however, unwarranted at hadron colliders, such as FCC-hh, where at least 100 signal events are typically required to suppress backgrounds in searches for rare prompt decays of Higgs and Z bosons [65, 66]. Future detector and experimental analysis progresses at the FCC-hh could make it less critical than for the LHC. It is conceivable that much more advanced data analysis techniques will be available by the time this new energy-frontier collider starts to operate, which can render observable some producible decays.

### v4

@@ II.1.3 Non-Relativistic QCD (NRQCD)
The formation of heavy quarkonium in the decay of a massive particle X is a multi-scale problem involving the hard scale given by $m_{\mathrm{X}}$, the heavy-flavour quark mass $m_{\mathrm{Q}}$, the relative momentum of the heavy quark pair $m_{\mathrm{Q}}v$, and the binding energy of the heavy quark pair $m_{\mathrm{Q}}v^{2}$, where $v$ is the typical relative velocity of the heavy quarks inside the meson. Since $m_{\mathrm{X}}\gg mv\gg mv^{2}\gg\Lambda_{\text{QCD}}$, the simple static limit applied for light quarks is not sufficient for these systems. Therefore, for heavy quark-antiquark bound states such as charmonium, bottomonium, and bottom-charm (Bc) mesons, a third class of models is often used where in addition to the usual expansion in powers of $\alpha_{\text{s}}$, the interactions are organized as an expansion in $v$. Since $v^{2}\approx 0.3,0.1$ for charmonium and bottomonium, respectively, the heavy quarks are nonrelativistic and described in a Non-Relativistic QCD (NRQCD) approach [Bodwin:1994jh]. In the NRQCD framework, the decay width of a heavy particle X into a given quarkonium meson M(n)(n) with quantum numbers $n=\,^{2s+1}L_{J}^{[1,8]}$ (corresponding to a state with spin $s$, orbital angular momentum $L$, total angular momentum $J$, in the $[1]$ or $[8]$ colour singlet or octet representations, respectively), $\mathrm{X}\to\mathrm{M}(n)+\mathrm{Y}$, factorizes as
@@
for the singlet charmonium ground states, where $N_{\mathrm{c}}=3$ is the number of colours. The LDMEs for the CS contributions can be obtained from the electromagnetic decays of quarkonia into dileptons or diphotons [Eichten:1995ch, Bodwin:2007fz], or from potential models. The NRQCD encapsulates two different mechanisms to describe the evolution of the heavy-flavour quark pair into a quarkonium meson, such as the colour-singlet model (CSM), and the colour-octet model (COM). In the CSM, the $\mathrm{Q}\overline{\mathrm{Q}}$ pairs are produced in CS states at the hard-scattering scale $m_{\mathrm{X}}$, and their quantum numbers are conserved during hadronization. The COM requires the extra radiation of gluons, and its contributions are more suppressed than in inclusive reactions, but has also been considered in a few cases. Beyond the non relativistic description of NRQCD, the relativistic quark model (RQM) [Ebert:2002pp] has been also employed in a few exclusive decay processes involving heavy quarkonia. The NRQCD calculations of heavy particle decays into quarkonia have been performed at LO, NLO, or next-to-NLO (NNLO) accuracy, resumming in some cases also leading (LL) or next-to-leading (NLL) logarithms.
As an example, the production of a bottom meson $\rm B\equiv(\mathrm{b}\overline{\mathrm{q}})_{n}=B_{\mathrm{u,d}},B_{\mathrm{s}},B_{\mathrm{c}}$ in the heavy particle decay process $\rm X\to\overline{\rm B}+Y$, can be described in a NRQCD-inspired heavy-quark recombination model [Braaten:2001bf] where the width takes the form
@@
with $(\mathrm{b}\overline{\mathrm{q}})_{n}$ representing the Fock state of the bottom b and accompanying $\rm\overline{q}$ quark, and the factor $\rho[(\mathrm{b}\overline{\mathrm{q}})_{n}\to\overline{\rm B}]$ is a nonperturbative probability for $(\mathrm{b}\overline{\mathrm{q}})_{n}$ to evolve into the $\rm\overline{\rm B}$ meson. In the case of $\rm B_{c}^{\pm}$ production, the nonperturbative transition $\rm(\mathrm{b}\overline{\mathrm{q}})_{n}\to\overline{\rm B}_{c}$ follows a definite velocity power-counting because the relative velocity of the charm and bottom quarks in the rest frame of the bottom meson (like that of the two bottom quarks in the $\Upsilon$ case) are small. This latter assumption is not fulfilled in the $\rm B_{\mathrm{u,d,s}}$ cases, where the Fock state $(\mathrm{b}\overline{\mathrm{q}})_{n}$ contributions to the $\rm\overline{\rm B}$ meson with different quantum numbers (e.g., in colour or angular momentum) are not necessarily suppressed, as aforementioned (although for practical purposes, the only relevant states are $n=\!^{1}S_{0}^{[1]},^{3}S_{1}^{[1]},^{1}S_{0}^{[8]},^{3}S_{1}^{[8]}$ [Braaten:2001bf]). By using the heavy-quark spin symmetry [Bodwin:1994jh], one can further reduce the nonperturbative transition probabilities $\rho$ from four to two as [dEnterria:2020ygk]
@@
where $m_{\mathrm{b}}$ is the bottom quark mass, $m_{\mathrm{{\overline{q}}}}=m_{\rm\bar{d}},m_{\rm\bar{s}}\approx 0.3$ GeV are the constituent light quark masses. The concrete application of the formulas above for the semiexclusive decay of a top quark into a B-meson plus an up-type quark, $\rm t\to\overline{\rm B}_{(s)}^{0}+q$ (discussed in Section VI.3), results in the following widths at LO in the QCD coupling $\alpha_{\mathrm{s}}$ [dEnterria:2020ygk]:
@@
with similar (yet longer) expressions available in Ref. [dEnterria:2020ygk] for the $\rm t\to\Upsilon+q$ decays.
@@ II.2 Experimental limits
As we will see below, predictions for rare decays of the electroweak and Higgs bosons, and the top quark, have branching fractions in the $10^{-5}$ to $10^{-15}$ range (or even down to $10^{-22}$ for positronium + photon final states). Such tiny branching fractions are very challenging experimentally, and no such decays have yet been observed for any of the particles. To provide an idea of the size of the data samples of W±, Z, H, and t particles discussed in this work, Table 1 collects their total number produced in past, current, and future colliders. The LEP numbers of W and Z bosons are obtained from the LEP-I and LEP-II electroweak summaries [ALEPH:2005ab, ALEPH:2013dgf], and indicate that no rare decay mode with rates below $10^{-7}$ and $10^{-5}$ for the Z and W bosons, respectively, has ever been probed in the clean experimental conditions of an $\mathrm{e}^{+}\mathrm{e}^{-}$ collider. We note, somehow curiously, that the last LEP-II operation, which integrated 2.46 fb-1 over $\sqrt{s}=189$–209 GeV [LEPWorkingGroupforHiggsbosonsearches:2003ing], featured a Higgs cross section of $\sigma_{\mathrm{H}}\approx 3$ fb (adding Higgstrahlung and weak-boson-fusion production processes) [dEnterria:2017jmj], and therefore LEP-II did produce a few Higgs(125 GeV) boson counts (however, the Higgs searches at the time were optimized for the associated production of a $m_{\mathrm{H}}=115$ GeV scalar boson plus an onshell Z boson, for which the H(125 GeV) signal would not have been visible). For the next $\mathrm{e}^{+}\mathrm{e}^{-}$ machine, we focus on FCC-ee because it is the planned facility with the largest data samples expected to be collected [FCC:2023]. The FCC-ee numbers in Table 1 cover the time span of the baseline 15-year program with four interaction points and four dedicated runs under consideration: Z-pole, WW threshold, HZ Higgstrahlung, and $\mathrm{t}\overline{\mathrm{t}}$ threshold. The largest data sample of all will be for Z bosons, for which $6\cdot 10^{12}$ particles will be produced. The number of H bosons at FCC-ee includes those produced in the HZ ($1.45\times 10^{6}$) and $\mathrm{t}\overline{\mathrm{t}}$ (+330+330k events) runs, plus via $\rm WW\to H$ at all runs (+125+125k). Although, in much more complicated background conditions than in $\mathrm{e}^{+}\mathrm{e}^{-}$ collisions, the HL-LHC (with 3 ab-1of p-p collisions at $\sqrt{s}=14$ TeV collected per ATLAS/CMS experiment) and the FCC-hh (p-p at $\sqrt{s}=100$ TeV with 30 ab-1) will truly serve as W, Z, Higgs, and top factories. We also list the number of the four massive particles produced at the Tevatron in $\mathrm{p}\mbox{-}\overline{\mathrm{p}}$ collisions at $\sqrt{s}=1.96$ TeV, as there exist still a couple of competitive rare decay limits from the CDF experiment. The hadron collider numbers in Table 1 are obtained from the corresponding production cross sections for each particle multiplied by the integrated luminosity quoted (plus the contribution from $\rm\mathrm{t}\overline{\mathrm{t}}\to W^{+}W^{-}+X$ decays for the W± counting). The cross sections at hadron colliders have been either obtained from the existing literature [Mangano:2016jyj] or, when not readily available or not fully up-to-date, have been recomputed at next-to-next-to-leading-order (NNLO) accuracy with mcfm v.8.0 [Boughezal:2016wmq] with the NNPDF3.1_NNLO parton distribution functions (PDFs) [NNPDF:2017mvq]. Theoretical (scale, PDF) uncertainties are at most 10% and not quoted.
@@
Table 1: Total number of Higgs, Z, and W bosons, and top quarks produced (or expected to be produced) in $\mathrm{e}^{+}\mathrm{e}^{-}$ collisions at LEP and FCC-$e$, as well as in $\mathrm{p}\mbox{-}\overline{\mathrm{p}}$ at Tevatron, and in p-p collisions at HL-LHC, and FCC-hh. For $\mathrm{e}^{+}\mathrm{e}^{-}$ colliders, the Z-pole, Higgstrahlung and Higgs weak-fusion, and pair production ($W$, $\mathrm{t}\overline{\mathrm{t}}$) cross sections (without ISR) are indicated. At LEP and FCC-$e$, the numbers of W bosons and top-quarks consider two bosons and two quarks produced per collision around the $\mathrm{e}^{+}\mathrm{e}^{-}\to\rm W^{+}W^{-},\,\mathrm{t}\overline{\mathrm{t}}$ thresholds. For hadron machines, the integrated luminosities and the NNLO production cross section for each particle are indicated. All top quark numbers are for pair production (i.e., the number of events are multiplied by two), and the number of W bosons is the (W+ + W-) sum including also the contributions from $\mathrm{t}\overline{\mathrm{t}}$ decays.
@@
$\sigma$(W) NN(W) $\sigma$(Z) NN(Z) $\sigma$(H) NN(H) $\sigma$($\mathrm{t}\overline{\mathrm{t}}$) NN(top)
@@
FCC-$e$ 2–10 pb $5\times 10^{8}$ 59 nb $6\times 10^{12}$ 200, 30 fb $1.9\times 10^{6}$ 0.5 pb $3.8\times 10^{6}$
@@
The largest yield increases across colliders are factors of $6\cdot 10^{3}$, $3\cdot 10^{5}$, and $4\cdot 10^{5}$ for W, Z, and H bosons, respectively, going from LEP to FCC-ee, and a factor of $9\cdot 10^{4}$ for top quarks from Tevatron to HL-LHC. As of today (end of 2024), measurements at LEP, Tevatron, and LHC have been able to set upper limits at 95% confidence level (CL) in about one-fourth of the $\sim$200 rare channels considered here. The most stringent experimental rare decays limits are listed in the tables below, including in some cases new results which are not yet available in the 2024 PDG decays listings [ParticleDataGroup:2024cfk]. One key goal of our work is to provide reasonable expectations of the achievable upper bounds at the HL-LHC, FCC-ee, and FCC-hh. The corresponding extrapolations are obtained through three different means:
@@
    For a few decay channels, there exist dedicated ATLAS and/or CMS studies that have determined the expected limits at the end of the HL-LHC phase [TheATLAScollaboration:2013nbo, ATLAS:2015xkp, CMS:2022kdd, Cerri:2018ypt, Liu:2020bem, ATLAS:2016qxw]. For such cases, we provide directly the expected limits (divided by $\sqrt{2}$ to account for the statistical combination of two experiments, ATLAS + CMS) with their bibliographical reference.
@@
    For those channels where LHC limits exist today in measurements with a given integrated luminosity (labelled e.g., $\mathcal{L}_{\mathrm{int}}$(13 TeV)” next), but for which no dedicated studies exist for HL-LHC, we estimate the latter by assuming that they will be statistically improved by the size of the final data sample, namely by the ratio of squared-root integrated luminosities $\sqrt{\smash{2\times 3~\mathrm{ab}^{-1}}/\smash{\mathcal{L}_{\mathrm{int}}(\text{13\,TeV})}}$, where the factor of two assumes an ATLAS + CMS combination. For individual LHC measurements carried out with the full Run-2 integrated luminosity of around 140 fb-1, this translates into an expected HL-LHC improvement of more than a factor of six. Bounds estimated that way are conservative, as they ignore improvements in the data analyses (e.g., optimization of the event selection, enlarged categorization depending on the production mode, adoption of more advanced statistical profiling methods, etc.), increased particle production cross sections (from collision energies rising from 13 to 14 TeV), and possible multi-experiment combinations (e.g., adding results from LHCb) for some channels. We have checked that our simple statistical approximation here gives limits comparable to those obtained with the method 1) above whenever dedicated studies are available.
@@
Finally, for the FCC-ee and FCC-hh cases, whenever a given decay channel has a branching fraction commensurate with the expected number of particles produced (i.e., whenever $\mathcal{B}\times N_{\mathrm{X}}\gtrsim 1$, where $N_{\mathrm{X}}$ are the FCC numbers given in Table 1 for any given particle), it will appear listed as “producible at FCC-ee/FCC-hh”. Assuming that potential backgrounds are small and/or well under control, as is often the case at $\mathrm{e}^{+}\mathrm{e}^{-}$ colliders, and even more so at FCC-ee where very precise detector performances are being required [FCC:2023], observation can be achieved with a handful of events. Such an assumption is, however, unwarranted at hadron colliders, such as FCC-hh, where at least 100 signal events are typically required to suppress backgrounds in searches for rare prompt decays of Higgs and Z bosons [CMS:2013rmy, ATLAS:2015rsn]. Future detector and experimental analysis progresses at the FCC-hh could make it less critical than for the LHC. It is conceivable that much more advanced data analysis techniques will be available by the time this new energy-frontier collider starts to operate, which can render observable some producible decays.